40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 19 18:40:00 2012, Jenne, Update, IOO, PZT1 removed, TT1 in place 
    Reply  Wed Dec 19 19:28:12 2012, Manasa, Update, IOO, PZT1 removed, TT1 in place 
       Reply  Thu Dec 20 10:11:12 2012, Manasa, Update, IOO, TT1 connections 
          Reply  Thu Dec 20 10:35:27 2012, Steve, Update, IOO, TT1 connections 
Message ID: 7857     Entry time: Wed Dec 19 18:40:00 2012     Reply to this: 7858
Author: Jenne 
Type: Update 
Category: IOO 
Subject: PZT1 removed, TT1 in place 

[Manasa, Jamie, Jenne]

PZT1 has been removed, and is wrapped in foil and stored in a (labeled) plastic box.

We beeped the cable between the cable holder bracket on the in-vac table, and the outside of the feedthrough.  Things are mirrored, so pins 1,14 (one edge on the feedthrough) go to pins 13,25 on the in-vac cable bracket.

Tip Tilt, serial number ### (Manasa will get the serial number and put it in the elog) was taken out of the cleanroom, for use as TT1.

We checked the epics controls from the TT screen that Jamie made a while back (accessible from the ASC tab on the sitemap) to the output of the AI board.  Things were very weird, but Jamie fixed them up in the model, then rebuilt and restarted the ASS model so that now the epics channel corresponding to, say, UL actually actuates on the UL output of the boards.

We tested the cables from the rack to the feedthrough, and discovered that they are also mirrored, to undo the mirroring between the feedthrough and the in-vac bracket.

Jamie made an adapter cable to take the pinout of the coil driver boards correctly to the pinout of the quadrupus cable, through this double-mirroring (i.e. no net mirror effect).

We set up a laser pointer on a tripod outside the door of the MC chamber (where the access connector usually is), and pointed it at the back of the TT.  Den or whomever put the cable on the TT didn't follow the diagram (or something got messed up somewhere), because when we actuate in pitch (+ on the uppers, - on the lowers), we see the TT move in yaw, and vice versa. 

We are in the process of removing the quadrupus from the TT, figuring out which connector goes where, putting it on correctly, and re-testing.

Depending on how far things get tonight, Jamie and Manasa may ask Steve to help them remove the BS door, so they can get started on replacing PZT2 with TT2.

ELOG V3.1.3-